a 250.0 ml buffer solution initially contains 0.15 M HCHO2 and 0.10 M NaCHO2. what is the ph of this buffer?
what component of the buffer must be added to change the ph to 3.91? what mass of this component must be added? assume negligible volume change
Answer #1

pH of buffer acidic bufferis given by, pH = pka + log(conjugate base/acid)

       pKa of Formic acid = 3.74

   pH = 3.74 +log(0.10/0.15)

      = 3.56

NaCHO2 component of the buffer must be added to change the pH from 3.56 to 3.91

Let x be mmoles of NaCHO2 added, then

   pH of buffer = pka + pka + log(conjugate base+x/acid)

            3.91= 3.74 + log((25+x)/37.5)

Number of mmoles of X(NaCHO2) must be added = 30.47

mass of X(NaCHO2) must be added = 30.47*68/1000 = 2.072 gms